A particle moves along the x-axis according to the expression x=3t² - 2t where x is in m and t is in s. What is the displacement of the particle in time intervals t=0 to t=1s and t=1s to t=4s


1
Expert's answer
2022-02-08T13:03:22-0500

Given:

"x=3t^2 - 2t"


(a)

"d_1=x(1)-x(0)=3*1^2-2*1-0=1\:\rm m"

(b)

"d_2=x(4)-x(1)\\
=(3*4^2-2*4)-(3*1^2-2*1)=39\:\rm m"